# Domain-adaptive Post-training of Large Language Models: An Empirical Study with Llama for Finance [Yifei: A more attractive/informative title?]

### v2

@@ 4 Experiments
Table 2: Results on similar (unseen) tasks. ‘*’ indicates that ‘GPT4o’ is used as the judge. Llama-Fin and its variant without PA (i.e., the ‘CPT+IT’ checkpoint) are highlighted in blue while the closed model is highlighted in gray. The best performing model for 8b on each benchmark is bolded. The overall best performance across all models is underlined. indicates that Llama-Fin outperforms the base Llama3-8b-inst.
@@
Table 3: Results on the novel tasks. The notations are the same as in Table 2. ‘Mcc’ refers to Matthews correlation coefficient, usually used in highly imbalanced data (Xie et al., 2024a).
We apply our method to the instruction-tuned Llama3-8b-inst, resulting in Llama-Fin. A summary of the hyper-parameters and computational resource requirements is given in Table F.1. For evaluation, we compare Llama-Fin with a wide range of baselines models, including its base model, Llama3-8b-inst, and the 8B peer, Llama3.1-8b-inst. We also include comparisons with models of other sizes, such as Phi-3.5-mini-instruct (Abdin et al., 2024) (3.8B), and Mistral-Nemo-inst (Jiang et al., 2023) (12B), as well as the closed model GPT-4o (OpenAI, 2023). Furthermore, we evaluated against the latest SoTA finance-specific LLM, Palmyra-Fin (70B) (Writer, 2024). Note that there are other financial LLMs available, such as FinMa (Xie et al., 2023), Finance-LLM (Cheng et al., 2024) and FinLLaVA (Xie et al., 2024a) that focuses on multi-modality. However, they are either not publicly available (FinLLaVA) or based on less advanced LLMs (e.g., LLaMA2). In our preliminary experiments, these models performed considerably worse than both our model and the baselines. Therefore, we have only included the SoTA financial LLM in our comparisons.
@@ 4.1 Main Results
Similar (unseen) tasks
To validate our approach, we first evaluate Llama-Fin on tasks that are similar (yet unseen) to the tasks used for training (e.g., test task EDTSUM (abstractive summarization) is similar to the training task TradeTheEvent (abstractive summarization)). From Table 2, we observe that Llama-Fin outperforms all other baselines in its size category by 10% - 25% absolute gain. It also surpasses significantly larger models, such as the finance-specific Palmyra-Fin-32K (70B). Notably, Llama-Fin also exceeds the performance of GPT-4o. These results are not very surprising since the test tasks are not entirely novel, but it demonstrates the effectiveness of our data and model recipe for domain-adaptive post-training.
Novel tasks
We now evaluate the generalization of Llama-Fin on the completely novel tasks that are also aligned to the expected capabilities (FinCap). Table 3 presents the results. Below, we summarize the key takeaways from the comparison:
@@ 4.2 Further Analysis and Ablations
Table 4 presents the effectiveness of PA on similar tasks. We see that PA leads to improved performance in 3 out of 5 tasks , while not causing any significant forgetting on the other two. Considering that these tasks do not need much reasoning, this is understandable as PA primarily targets the reasoning tasks.
In Table 5, we show the same ablation for the novel tasks. In Concept (rows 9-12) and IF/chat (row 30) capabilities, removing PA often leads to worse results, indicating its effectiveness. In Task (rows 13-29), we see a mixed performance with and without PA. This is again not surprising as PA specifically focuses on reasoning tasks. Interestingly, we observe that for certain tasks (e.g., CRA-TravelInsurance, CRA-Taiwan and CRA-ProroSeguro), PA negatively impacts performance, resulting in worse outcomes compared to without PA. Even GPT-4o performs poorly in these tasks. This suggests that for some tasks, leveraging reasoning capabilities might not be beneficial, as these tasks could be inherently ‘easy’ and solvable without the need for explicit reasoning. Such observations align with prior findings (Sprague et al., 2024; Liu et al., 2024). In reasoning tasks (rows 31-38), Llama-Fin is significantly better than without PA variant, further confirming that our proposed FAP and SCP are particularly effective in improving reasoning performance beyond the already strong checkpoint of Llama-Fin (w/o PA).
@@ 5 Conclusion
We introduce FinDaP, an open SoTA finance-specific post-training framework, consists of FinCap that identifies four key capabilities, FinRec which jointly trains CPT and IT, and constructing PA preference data with stepwise signals, FinTrain that implemnts FinRec and FinEval a comprehensive evaluation setups. Under FinDaP, we develop Llama-Fin, a SoTA finance LLM. In this development, we conduct a systematic study on effectively adapting a target domain through post-training. For each stage, we reveal the distinct challenges, objectives, and effective strategies. Looking ahead, we aim to explore additional domain-specific capabilities using FinDaP.
@@ 6 Limitations
Appendix A Our Contributions
In summary, our key contributions are:
  • A comprehensive guidance for finance-specific post-training, including identification of capabilities, evaluation, data and model recipe design.
  • A systematic exploration on each stage of post-training, with an emphasis on the goals, challenges and effective approaches at each stage.
  • A novel preference alignment approach that constructs preference data using on-policy trajectories guided by outcome and process signals.
  • The above yields a new state-of-the-art finance LLM (Llama-Fin) at the 8b parameter scale.
Appendix B Ablation and Understanding FinDaP
@@ B.2 Instruction Following
From Prompt to IT Data.
We introduced our prompt curation in §D. We create the responses for IT by filtering existing responses or creating new responses. For prompts with existing responses, we generally keep the original responses if they were written by a human or a strong model, such as GPT-4. We also filter out empty responses. For prompts without responses, for example, exercises extracted from books that may not have solutions provided, we generate new responses using GPT-4o. Similar to CPT data, we construct three versions of IT data, IT-In, which contains only financial (in-domain) tasks, IT-Gen, which contains only general tasks, and IT-Mix, which includes a mixture of the IT-In and IT-Gen.
@@ B.3 Combining CPT and IT
A natural choice is to conduct CPT and IT sequentially (Lambert et al., 2024). On the one hand, this is flexible as it allows for different settings (e.g., data size) in each stage. On the other hand, it does not help prevent forgetting during the CPT stage, leaving the LLM dependent on IT to ‘recover’ its instruction-following capability. To make a more grounded decision, we conduct experiments on both sequential and joint training approaches. In joint training, an additional hyperparameter to consider is the mixture ratio. We find that down-sampling CPT data to match the size of IT data is the most effective strategy.
@@ B.4 Preference Alignment
Figure B.7: Average performance on selected datasets for training from ‘CPT+IT’ on PA using Trajectory-GenORM, noted as PA (GenORM) and Trajectory-GenORM + Trajectory-GenPRM, noted as PA (GenORM+GenPRM).
@@ Appendix D FinTrain
Capability Domain CPT Dataset Size Reference
@@
Capability Domain Task IT Dataset Size Reference
@@
Prompt for Filtering the Text Below is an extract from a text book. Evaluate whether the book has a high financial value and could be useful in an financial setting for teaching financial students using the additive scoring system described below. Points are accumulated strictly based on the satisfaction of each criterion: - Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take finacial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take financial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided. Add another point if the extract addresses certain elements pertinent to finance and aligns closely with financial standards. It might offer a superficial overview of potentially useful topics or present information in a disorganized manner and incoherent writing style. Award a third point if the extract is appropriate for financial use and introduces key concepts relevant to financial curricula. It is coherent and comprehensive. Grant a fourth point if the extract is highly relevant and beneficial for financial learning purposes for a level not higher than financial students, exhibiting a clear and consistent writing style. It offers substantial financial content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for financial students. The content is coherent, focused, and valuable for structured learning. Bestow a fifth point if the extract is outstanding in its financial value, perfectly suited for teaching either at financial students. It follows detailed reasoning, the writing style is easy to follow and offers profound and thorough insights into the subject matter, devoid of any non-financial or complex content. The extract: <EXAMPLE>. After examining the extract, You will output a json object containing the following 2 fields: { "Justification": string // Briefly justify your total score, up to 100 words. "Score": integer // Conclude with the score }
@@
Figure D.2 shows the prompt we used to extract exercises from books. We carefully design the prompt to extract both the question part of an exercise, which potentially include questions, scenario and exhibits, and the answer part of the exercise, which may include answer choices and solution. In books, questions and their corresponding answers can be located far apart (e.g., the questions may appear at the beginning while the solutions are provided at the end), meaning they may not be captured within the same chunk. As a result, some questions may not have corresponding extracted answers. For such cases, GPT-4o’s generated answers are used when converting the prompt into instruction-following or preference-alignment data, as detailed in Sections B.2 and 3.2.1.
@@ Appendix E FinEval
Appendix F Summary of the Final Recipe and Hyper-parameters
@@
Attention CPT: full attention with cross-docuemnt attention masking
IT: full attention with instruction mask-out and cross-docuemnt attention masking
@@
Data Trajectory-GenORM and Trajectory-GenPRM
@@
Table F.1: Final recipe of Llama-Fin. The joint training of CPT and IT is structured into two groups, with each group undergoing joint training sequentially. The second group utilizes higher-quality data (sourced from books), following the typical curriculum training practice (Gao et al., 2024). For PA, we employ a modified DPO loss with an additional negative log-likelihood term, similar to Pang et al. (2024), as it has shown to be more effective than relying solely on the original DPO loss.
Appendix G GenRM Details
In §3.2.1, we simplified the prompt for GenRM for the purpose of illustration in Figure 2. In this section, we give full detailed of the prompt for Final Answer Preference (FAP) and Stepwise Corrective Preference (SCP) in Figure G.1 and Figure G.2, respectively.

### v4

@@ 4 Experiments
Table 2: Results on similar (unseen) tasks. Llama-Fin is highlighted in blue while the closed model is highlighted in gray. The best performing model for 8b on each benchmark is bolded. The overall best performance across all models is underlined. indicates that Llama-Fin outperforms the base Llama3-8b-inst.
@@
Table 3: Results on the novel tasks. The notations are the same as in Table 2. ‘*’ indicates that ‘GPT4o’ is used as the judge. ‘Mcc’ refers to Matthews correlation coefficient, usually used in highly imbalanced data (Xie et al., 2024a).
We apply our method to the instruction-tuned Llama3-8b-inst, resulting in Llama-Fin (GPT-4o is used as GenRM). A summary of the hyper-parameters and computational resource requirements is given in Table G.1. For evaluation, we compare Llama-Fin with a wide range of baselines models, including its base model, Llama3-8b-inst, and the 8B peer, Llama3.1-8b-inst. We also include comparisons with models of other sizes, such as Phi-3.5-mini-instruct (Abdin et al., 2024) (3.8B), and Mistral-Nemo-inst (Jiang et al., 2023) (12B), as well as the closed model GPT-4o (OpenAI, 2023). Furthermore, we evaluated against the latest SoTA finance-specific LLM, Palmyra-Fin (70B) (Writer, 2024). Note that there are other financial LLMs available, such as FinMa (Xie et al., 2023) and FinLLaVA (Xie et al., 2024a). However, they are either not publicly available (FinLLaVA) or based on less advanced LLMs (e.g., LLaMA2). In preliminary experiments, these models performed considerably worse than our model (see App. F). Therefore, we have only included the SoTA financial LLM in our comparisons.
@@ 4.1 Main Results
Similar (unseen) tasks. To validate our approach, we first evaluate Llama-Fin on tasks that are similar (yet unseen) to the tasks used for training (e.g., test task EDTSUM (abstractive summarization) is similar to the training task TradeTheEvent (abstractive summarization)). From Table 2, we observe that Llama-Fin outperforms all other baselines in its size category by 10% - 25% absolute gain. It also surpasses significantly larger models, such as the finance-specific Palmyra-Fin (70B). Notably, Llama-Fin also exceeds the performance of GPT-4o. These results are not very surprising since the test tasks are not entirely novel, but it demonstrates the effectiveness of our data and model recipe for domain-adaptive post-training.
Novel tasks. We now evaluate the generalization of Llama-Fin on the completely novel tasks that are also aligned to the expected capabilities (FinCap). Table 3 presents the results. Below, we summarize the key takeaways from the comparison:
@@ 4.2 Further Analysis and Ablations
Table 4 presents the effectiveness of PA on similar tasks. We see that PA leads to improved performance in 3 out of 5 tasks, while not causing any significant forgetting on the other two. This is expected as PA primarily targets the reasoning tasks whereas these tasks do not need much reasoning.
In Table 5, we show the same ablation for the novel tasks. In Concept (rows 2-5) and IF/chat (row 23) capabilities, removing PA often leads to worse results, indicating its effectiveness. In Task (rows 6-22), we see a mixed performance with and without PA. This is again not surprising as PA specifically focuses on reasoning tasks. Interestingly, we observe that for certain tasks (e.g., CRA-TravelInsurance, CRA-Taiwan and CRA-ProroSeguro), PA negatively impacts performance, resulting in worse outcomes compared to without PA. Even GPT-4o performs poorly in these tasks. This suggests that for some tasks, leveraging reasoning capabilities might not be beneficial, as these tasks could be inherently “easy” and solvable without the need for explicit reasoning. Such observations align with prior findings (Sprague et al., 2024; Liu et al., 2024). In Reasoning (rows 24-31), Llama-Fin is significantly better than without PA variant, further confirming that our proposed FAP and SCP are particularly effective in improving reasoning performance beyond the already strong checkpoint of Llama-Fin (w/o PA).
@@ 5 Conclusion
We introduce FinDaP, an open SoTA finance-specific post-training framework, consists of FinCap that identifies four key capabilities; FinRec which jointly trains CPT and IT, and constructing PA preference data with stepwise signals; FinTrain that implements FinRec; and FinEval, a comprehensive evaluation setup. Under FinDaP, we develop Llama-Fin, a SoTA finance LLM. In this development, we conduct a systematic study on effectively adapting a target domain through post-training. For each stage, we reveal the distinct challenges, objectives, and effective strategies. Looking ahead, we aim to scale up the base LLM and explore additional domain-specific capabilities using FinDaP.
@@ 6 Limitations
Appendix A Preventing Data Contamination
When designing FinEval, we took extra care to ensure that evaluation tasks do not duplicate any samples from FinTrain. To further verify this, we followed your suggestion and computed string matches between FinEval and FinTrain.
Specifically, we adopted the decontamination procedure described in the Hugging Face blog you referenced. A training sample is contaminated if it is overlapped with any evaluation sample. The contamination ratio is computed as the fraction of contamination samples in the training samples. Based on this criterion, we report two contamination ratios:
  • 0% under the strictest setting, where only complete sample matches are considered contamination.
  • 0.003% using the method described in the blog—where 10-gram matches are used for pre-identification, followed by difflib.SequenceMatcher. If over 50% of its characters match any of the evaluation samples, the training sample is considered contaminated.
These contamination rates are extremely low, indicating minimal overlap between FinTrain and FinEval. Upon manual inspection of the few samples flagged by the 50% character overlap rule, we found they involve either (1) partial overlap in the question format or instruction prompt, which is expected for the similar tasks where the task type (e.g., sentiment analysis) has been seen, but the content remain unseen; or (2) partial overlap in the input content (e.g., shared elements in bank transcripts), but the specific question and answers are unseen. In both cases, these do not indicate memorization or leakage of benchmark content.
Appendix B Ablations and Understanding FinDaP
@@ B.2 Instruction Following
From Prompt to IT Data. We introduced our prompt curation in §D. We create the responses for IT by filtering existing responses or creating new responses. For prompts with existing responses, we generally keep the original responses if they were written by a human or a strong model, such as GPT-4. We also filter out empty responses. For prompts without responses, for example, exercises extracted from books that may not have solutions provided, we generate new responses using GPT-4o. Similar to CPT data, we construct three versions of IT data, IT-In, which contains only financial (in-domain) tasks, IT-Gen, which contains only general tasks, and IT-Mix, which includes a mixture of the IT-In and IT-Gen.
@@ B.3 Combining CPT and IT
A natural choice is to conduct CPT and IT sequentially (Lambert et al., 2024). On the one hand, this is flexible as it allows for different settings (e.g., data size) in each stage. On the other hand, it does not help prevent forgetting during the CPT stage, leaving the LLM dependent on IT to ‘recover’ its instruction-following capability. To make a more grounded decision, we conduct experiments on both sequential and joint training approaches. In joint training, an additional hyperparameter to consider is the mixture ratio. We down-sample CPT data to match the size of IT data. In “Other sampling strategies” section, we will show this is the most effective strategy.
@@ Appendix D FinTrain
Capability Domain Dataset Size Reference
@@
Capability Domain Task Dataset Size Reference
@@
Prompt for Filtering the Text Below is an extract from a text book. Evaluate whether the book has a high financial value and could be useful in an financial setting for teaching financial students using the additive scoring system described below. Points are accumulated strictly based on the satisfaction of each criterion: Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take financial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided. Add another point if the extract addresses certain elements pertinent to finance and aligns closely with financial standards. It might offer a superficial overview of potentially useful topics or present information in a disorganized manner and incoherent writing style. Award a third point if the extract is appropriate for financial use and introduces key concepts relevant to financial curricula. It is coherent and comprehensive. Grant a fourth point if the extract is highly relevant and beneficial for financial learning purposes for a level not higher than financial students, exhibiting a clear and consistent writing style. It offers substantial financial content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for financial students. The content is coherent, focused, and valuable for structured learning. Bestow a fifth point if the extract is outstanding in its financial value, perfectly suited for teaching either at financial students. It follows detailed reasoning, the writing style is easy to follow and offers profound and thorough insights into the subject matter, devoid of any non-financial or complex content. The extract: <EXAMPLE>. After examining the extract, You will output a json object containing the following 2 fields: { "Justification": string // Briefly justify your total score, up to 100 words. "Score": integer // Conclude with the score }
@@
Figure D.2 shows the prompt we used to extract exercises from books. We carefully design the prompt to extract both the question part of an exercise, which potentially include questions, scenario and exhibits, and the answer part of the exercise, which may include answer choices and solution. In books, questions and their corresponding answers can be located far apart (e.g., the questions may appear at the beginning while the solutions are provided at the end), meaning they may not be captured within the same chunk. As a result, some questions may not have corresponding extracted answers. For such cases, GPT-4o’s generated answers are used when converting the prompt into instruction-following or preference-alignment data.
@@ Appendix E FinEval
Appendix F Preliminary Experiments on Older Financial LLMs
As mentioned in Section 4, the reason we did not include older financial LLMs is that they are either not publicly available (e.g., Bloomberg GPT) or clearly worse than our model. As a result, we only include the SoTA finance LLM (i.e., Palmyra Fin 70b) in the comparison.
To further support this point, we compare performance on overlapping evaluation benchmarks, using the reported numbers for other baselines extracted from their papers. We made careful efforts to ensure comparability:
$\bullet$ Metrics. We noticed that different metrics were used across baselines and our methods. For example, some baselines reported F1 scores for FPB and FiQA SA, while we originally reported accuracy. For NER, the baselines used Entity F1, whereas we initially reported ROUGE scores. To ensure fair comparison, we re-ran our evaluation using the same metrics. We reported both accuracy and F1 for FPB and FiQA SA, and used Entity F1 for NER.
$\bullet$ Test Datasets. The test datasets are the same. We follow the datasets used in Xie et al. (2023) [^6] https://huggingface.co/collections/TheFinAI/english-evaluation-dataset-658f515911f68f12ea193194, which include 235 test samples for FiQA SA, 970 samples for FPB, and 98 samples for NER. These statistics also match those reported in Table E.1 of our Appendix. We do not use the training or validation sets, as our evaluation is conducted purely in the zero-shot setting. The baseline results are taken directly from Table 5 in Xie et al. (2023), which ensures consistency in comparison and also corresponds to Table 1 in Xie et al. (2024b).
Table F.1 shows the results. These results clearly show that our model outperforms these older financial LLMs, including significantly larger models such as FinMA 30B. Moreover, their reported results are based on few-shot settings, whereas our evaluations are conducted in the zero-shot setting, further highlighting the effectiveness of our approach.
Dataset Metric
Llama-Fin
8b
Bloomberg
GPT
FinPythia
7B
FinMA
7B
FinMA
30B
FPB Acc 91.13 59.90 86.00 87.00
F1 91.28 51.07 64.43 86.00 88.00
FiQA SA Acc 95.32 52.34 84.00 87.00
F1 95.39 75.07 53.04
NER EntityF1 77.09 60.82 48.42 75.00 62.00
Table F.1: Experiments on older baselines.
Appendix G Summary of the Final Recipe and Hyper-parameters
@@
Attention CPT: full attention with cross-document attention masking
IT: attention with instruction mask-out and cross-document attention masking
@@
Data FAP and SCP
@@
Attention Attention with instruction mask-out and cross-document attention masking
@@
Table G.1: Final recipe of Llama-Fin. The joint training of CPT and IT is structured into two groups, with each group undergoing joint training sequentially. The second group utilizes higher-quality data (sourced from books), following the typical curriculum training practice (Gao et al., 2024). For PA, we employ a modified DPO loss with an additional negative log-likelihood term, similar to Pang et al. (2024), as it has shown to be more effective than relying solely on the original DPO loss.
Appendix H GenRM Prompt Details
In Figure 2, we simplified the prompt for GenRM for the purpose of illustration. In this section, we give full detailed of the prompt for Final Answer Preference (FAP) and Stepwise Corrective Preference (SCP) in Figure H.1 and Figure H.2, respectively.